# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 4. Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
When $\min\{|I|,|J|\}<h-1$, the proof is easier. If $|I|,|J|<h-1$, then we have two patterns: either there are two visits to 00 and two visits to the same other position, or there are three visits to 00. In either case, we only need to sum over $x_{i}$ or $y_{j}$ for at most $h-3$ times . Hence, we do not need to introduce the factor $(1+|x_{m}-x_{n}|^{2a})/(1+|y_{k}-y_{l}|^{2a})$. If $|I|<h-1$ and $|J|=h-1$, then we can find $k,l$ belonging to the same partition element of $I$ with $k,j\notin J(1)$. Then we can replace the factor $(1+|x_{m}-x_{n}|^{2a})/(1+|y_{k}-y_{l}|^{2a})$ in (4.34) by $1/(1+|y_{k}-y_{l}|^{2a})$ (and also its reciprocal). The rest of the proof is then the same as that for $|I|=|J|=h-1$.
Then we turn to (4.32). The treatment is similar as above. If $|I|=|J|=h-1$, then we have Case (1) and Case (2) as above. We will insert different factors according to $k,l,m,n$ and $K$:
@@
The reason behind this strategy is that when $j\notin K$, $(1+|y_{j}-x_{j}|^{2})^{-1}$ is summable and we save the degree of power by $\frac{1}{2}$.
@@
We only treat $Q_{\lambda,N}^{I,*}$, since the treatment for $Q_{\lambda,N}^{*,I}$ is the same. It suffices to prove that there exits come constant $c\in(0,+\infty)$, such that for $h\geq 3$ and all $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{q}(({\mathbbm{Z}})^{h}_{J})$,
(4.38) $\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}}\varphi(\boldsymbol{x})Q^{I,*}_{\lambda,N}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq cN^{\frac{1}{p}}\|\varphi\|_{\ell^{p}}\|\psi\big{\|}_{\ell^{q}}$
uniformly for all $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{p}(({\mathbbm{Z}})^{h}_{J})$.
@@
Then for any give $K\subsetneqq\{1,\cdots,h\}$, we are going to bound
(4.39) $\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}}f(\boldsymbol{x})\bigg{(}\sum\limits_{n=1}^{2N}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{2})\sqrt{n}}\bigg{)}g(\boldsymbol{y}),$
which can also be done by splitting the cases $|\boldsymbol{x}-\boldsymbol{y}|\leq\sqrt{N}$ and $|\boldsymbol{x}-\boldsymbol{y}|\geq\sqrt{N}$. For $|\boldsymbol{x}-\boldsymbol{y}|\leq\sqrt{N}$, if $1,2\in K$, then we insert a factor $\log(1+\frac{N}{y_{1}^{2}})$ and its reciprocal. If either $1$ or $2$ does not belong to $K$, then we do not need to insert any factor.
@@
It suffices to prove that there exists some constant $c\in(0,+\infty)$, such that
(4.40) $\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{h}}\varphi(\boldsymbol{x})U_{\lambda,N,\widetilde{N}}^{I}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\frac{c}{\log\big{(}\widehat{\lambda}\frac{\widetilde{N}}{N}\big{)}\sigma_{\widetilde{N}}^{2}}\|\varphi\|_{\ell^{p}}\|\psi\big{\|}_{\ell^{q}}$
uniformly for all $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{p}(({\mathbbm{Z}})^{h}_{J})$.
Recall (4.15) and (4.13) and we assume that $I(1)=\{1,2\}$, that is, only the random walks with indices 1 and 2 visits 00. We have that
@@
By (2.26) and setting $T:=\frac{\widetilde{N}}{N}\geq 1$, we have that
$\begin{split}\sum\limits_{\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{h}}U_{\frac{\widehat{\lambda}}{N},N,\widetilde{N}}^{I}(\boldsymbol{x},\boldsymbol{y})\leq 1+C\log\widetilde{N}\sum\limits_{n=1}^{2\widetilde{N}/T}e^{-\widehat{\lambda}\frac{nT}{\widetilde{N}}}G_{\vartheta}\Big{(}\frac{n}{\widetilde{N}}\Big{)}\frac{1}{\widetilde{N}}\leq 1+C\log\widetilde{N}\int_{0}^{\frac{2}{T}}e^{-\widehat{\lambda}Tt}G_{\vartheta}(t){\mathrm{d}}t.\end{split}$
By (2.28), $G_{\vartheta}(t)$ is uniformly bounded on $[\frac{1}{2},2]$ and $G_{\vartheta}(t)\sim[t(\log\frac{1}{t})^{2}]^{-1}$ as $t\to 0$. We obtain that
$\begin{split}\int_{0}^{\frac{2}{T}}e^{-\widehat{\lambda}Tt}G_{\vartheta}(t){\mathrm{d}}t&\leq C\int_{\frac{1}{2}}^{2}e^{-\widehat{\lambda}Tt}{\mathrm{d}}t+\int_{0}^{(\widehat{\lambda}T)^{-\frac{1}{2}}}\frac{{\mathrm{d}}t}{t(\log\frac{1}{t})^{2}}+e^{-\sqrt{\widehat{\lambda}T}}\int_{(\widehat{\lambda}T)^{-\frac{1}{2}}}^{\frac{1}{2}\wedge\frac{2}{T}}\frac{{\mathrm{d}}t}{t(\log\frac{1}{t})^{2}}\leq\frac{C}{\log\widehat{\lambda}T},\end{split}$
@@
5. Approximating ${\mathcal{Z}}_{N}(\varphi,\psi)$ by the coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
@@
5.1. Approximating ${\mathcal{Z}}_{N}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$
@@
(5.1) $\begin{split}&\Big{\|}\Big{(}{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}-{\mathcal{Z}}_{N}\Big{)}(\varphi,\psi)\Big{\|}_{2}^{2}=\Big{(}\sum\limits_{{\mathcal{C}}_{1}}+\sum\limits_{{\mathcal{C}}_{2}}\Big{)}\frac{1}{N}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\\ &\left(q_{0,d_{1}}^{N}(\varphi,0)\right)^{2}\overline{U}_{N}(f_{1}-d_{1})\Big{\{}\prod\limits_{j=2}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\Big{\}}\left(q_{f_{k},N}^{N}(0,\psi)\right)^{2},\end{split}$
@@
(5.2) $\limsup\limits_{N\to\infty}I_{N,\varepsilon}\leq\frac{C}{\log\frac{1}{\varepsilon}}\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2},$
(5.3) $\limsup\limits_{N\to\infty}J_{N,\varepsilon}\leq C\bigg{(}\int_{0}^{K_{\varepsilon}\varepsilon}{\mathrm{d}}s\int_{\mathbbm{R}}g_{s}^{2}(x){\mathrm{d}}x+\frac{(\log K_{\varepsilon})^{2}}{\log\frac{1}{\varepsilon}}\bigg{)}\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2}.$
@@
(5.4) $\lim\limits_{N\to\infty}\frac{1}{N}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}_{k})}\left(q_{0,d_{1}}^{N}(\varphi,0)\right)^{2}\overline{U}_{N}(f_{1}-d_{1})\bigg{\{}\prod\limits_{j=2}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\bigg{\}}={\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})$
@@
(5.5) ${\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}}):=\idotsint\limits_{a_{1}\leq b_{1}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{1}),\cdots,a_{k}\leq b_{k}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{k})}g_{a_{1}}(\varphi,0)^{2}G_{\vartheta}(b_{1}-a_{1})\bigg{\{}\prod\limits_{j=2}^{k}\frac{G_{\vartheta}(b_{j}-a_{j})}{a_{j}-b_{j-1}}\bigg{\}}{\mathrm{d}}a_{1}{\mathrm{d}}b_{1}\cdots{\mathrm{d}}a_{k}{\mathrm{d}}b_{k},$
@@
Recall the local limit (2.20) for $u(n)$, the asymptotics (2.27) and the uniform bound (2.26) for $\overline{U}_{N}(n)$. The lemma then follows by a Riemann sum approximation, similar to the proof of Proposition 2.3. ${\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})$ is well-defined by the renewal property (6.14) in [CSZ19a]:
(5.6) $G_{\vartheta}(t)=\iint_{0<u<\bar{t}\leq v<t}G_{\vartheta}(u)\frac{1}{v-u}G_{\vartheta}(t-v){\mathrm{d}}u{\mathrm{d}}v,\quad\forall t\in(\bar{t},+\infty).$
@@
(5.7) ${\mathcal{I}}_{\varepsilon}^{s,t}({\mathbf{I}}):=\idotsint\limits_{s\leq b_{1}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{1}),\cdots,a_{k}\leq t\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{k})}G_{\vartheta}(b_{1}-s)\bigg{\{}\prod\limits_{j=2}^{k-1}\frac{G_{\vartheta}(b_{j}-a_{j})}{a_{j}-b_{j-1}}\bigg{\}}\frac{G_{\vartheta}(t-a_{k})}{a_{k}-b_{k-1}}{\mathrm{d}}b_{1}{\mathrm{d}}a_{2}\cdots{\mathrm{d}}b_{k-1}{\mathrm{d}}a_{k},$
@@
(5.8) $\sum\limits_{k=1}^{\infty}\sum\limits_{\mathrm{j}=:\mathrm{i}_{1}<\mathrm{i}_{2}<\cdots<\mathrm{i}_{k-1}<\mathrm{i}_{k}:=\mathrm{j}^{\prime}}{\mathcal{I}}_{\varepsilon}^{s,t}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=G_{\vartheta}(t-s)$
@@
(5.9) $\limsup\limits_{N\to\infty}I_{N,\varepsilon}$ $\leq\|\psi\|_{\infty}\sum\limits_{k\geq(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{0<\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}\leq\frac{1}{\varepsilon}}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=:I_{\varepsilon},$
(5.10) $\limsup\limits_{N\to\infty}J_{N,\varepsilon}$ $\leq\|\psi\|_{\infty}\sum\limits_{k=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in({\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}})^{c}}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=:J_{\varepsilon}.$
@@
$I_{\varepsilon}\leq\frac{1}{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{{\mathbf{I}}\subset\{1,\cdots,\frac{1}{\varepsilon}\},|{\mathbf{I}}|\geq(\log\frac{1}{\varepsilon})^{2}}|{\mathbf{I}}|{\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})\leq\frac{1}{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{j=1}^{\frac{1}{\varepsilon}}\sum\limits_{j\in{\mathbf{I}}\subset\{1,\cdots,\frac{1}{\varepsilon}\}}{\mathbf{I}}_{\varepsilon}^{\varphi}({\mathbf{I}}).$

### v3

@@ 4 Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
When $\min\{|I|,|J|\}<h-1$, the proof is easier. If |I|,|J|<h1|I|,|J|<h-1, then we have two patterns: either there are two random walks visiting 0 and two other random walks visiting the same other position, or there are three random walks visiting 0. In either case, we only need to sum over $x_{i}$ or $y_{j}$ for at most h3h-3 times . Hence, we do not need to introduce the factor $(1+|x_{m}-x_{n}|^{2a})/(1+|y_{k}-y_{l}|^{2a})$. If |I|<h1|I|<h-1 and |J|=h1|J|=h-1, then we can find $k,l$ belonging to the same partition element of $I$ with $k,j\notin J(1)$. Then we can replace the factor $(1+|x_{m}-x_{n}|^{2a})/(1+|y_{k}-y_{l}|^{2a})$ in (4.34) by $1/(1+|y_{k}-y_{l}|^{2a})$ (and also its reciprocal, resp.). The rest of the proof is then the same as that for |I|=|J|=h1|I|=|J|=h-1.
Then we turn to (4.32). The treatment is similar. If |I|=|J|=h1|I|=|J|=h-1, then we have Case (1) and Case (2) as above. We will insert different factors according to $k,l,m,n$ and $K$:
@@
The reason behind this strategy is that when $j\notin K$, $(1+|y_{j}-x_{j}|^{1+\gamma})^{-1}$ is summable and we save the degree of power by $\frac{1}{2}$.
@@
We only treat $Q_{\lambda,N}^{I,*}$, since the treatment for $Q_{\lambda,N}^{*,I}$ is the same. It suffices to prove that there exits come constant $c\in(0,+\infty)$, such that uniformly for all $N$ and $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{p}(({\mathbbm{Z}})^{h}_{J})$.
$\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}}\varphi(\boldsymbol{x})Q^{I,*}_{\lambda,N}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq cN^{\frac{1}{p}}\|\varphi\|_{\ell^{p}}\|\psi\big\|_{\ell^{q}}.$ (4.38)
@@
Then for any given $K\subsetneqq\{1,\cdots,h\}$, we are going to bound
$\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}}\varphi(\boldsymbol{x})\bigg(\sum\limits_{n=1}^{2N}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{1+\gamma})\sqrt{n}}\bigg)\psi(\boldsymbol{y}),$ (4.39)
which can also be done by splitting the cases $|\boldsymbol{x}-\boldsymbol{y}|\leq\sqrt{N}$ and $|\boldsymbol{x}-\boldsymbol{y}|\geq\sqrt{N}$, after we have applied Hölder’s inequality and summed over $x_{j},y_{j}$ for $j\notin K$. For $|\boldsymbol{x}-\boldsymbol{y}|\leq\sqrt{N}$, if $1,2\in K$, then we insert a factor $\log(1+\frac{N}{y_{1}^{2}})$ and its reciprocal. If either $1$ or $2$ does not belong to $K$, then we do not need to insert any factor. ∎
@@
It suffices to prove that there exists some constant $c\in(0,+\infty)$, such that uniformly for all $N,\widetilde{N}$ and $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{p}(({\mathbbm{Z}})^{h}_{J})$,
$\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{h}}\varphi(\boldsymbol{x})U_{\lambda,N,\widetilde{N}}^{I}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\frac{c}{\log\big(\widehat{\lambda}\frac{\widetilde{N}}{N}\big)\sigma_{\widetilde{N}}^{2}}\|\varphi\|_{\ell^{p}}\|\psi\big\|_{\ell^{q}}.$ (4.40)
Recall (4.15) and (4.13) and we assume that $I(1)=\{1,2\}$, that is, only the random walks with indices 1 and 2 visit 0. We have that
@@
Set $T:=\frac{\widetilde{N}}{N}\geq 1$. By (2.26), $\sum_{\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{h}}U_{\widehat{\lambda}/N,N,\widetilde{N}}^{I}(\boldsymbol{x},\boldsymbol{y})$ is bounded above by
$\begin{split}1+C\log\widetilde{N}\sum\limits_{n=1}^{2\widetilde{N}/T}e^{-\widehat{\lambda}\frac{nT}{\widetilde{N}}}G_{\vartheta}\Big(\frac{n}{\widetilde{N}}\Big)\frac{1}{\widetilde{N}}\leq 1+C\log\widetilde{N}\int_{0}^{\frac{2}{T}}e^{-\widehat{\lambda}Tt}G_{\vartheta}(t){\mathrm{d}}t.\end{split}$
By (2.28), $G_{\vartheta}(t)$ is uniformly bounded on $\frac{1}{2},2]$ and $G_{\vartheta}(t)\sim[t(\log\frac{1}{t})^{2}]^{-1}$ as $t\to 0$. We have that $\int_{0}^{2/T}e^{-\widehat{\lambda}Tt}G_{\vartheta}(t){\mathrm{d}}t$ is bounded above by
$\begin{split}C\int_{\frac{1}{2}}^{2}e^{-\widehat{\lambda}Tt}{\mathrm{d}}t+\int_{0}^{(\widehat{\lambda}T)^{-\frac{1}{2}}}\frac{{\mathrm{d}}t}{t(\log\frac{1}{t})^{2}}+e^{-\sqrt{\widehat{\lambda}T}}\int_{(\widehat{\lambda}T)^{-\frac{1}{2}}}^{\frac{1}{2}\wedge\frac{2}{T}}\frac{{\mathrm{d}}t}{t(\log\frac{1}{t})^{2}}\leq\frac{C}{\log\widehat{\lambda}T},\end{split}$
@@
5 Approximating ${\mathcal{Z}}_{N}(\varphi,\psi)$ by the coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
@@
5.1 Approximating ${\mathcal{Z}}_{N}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$
@@
$\begin{split}&\Big\|\Big({\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}-{\mathcal{Z}}_{N}\Big)(\varphi,\psi)\Big\|_{2}^{2}=\Big(\sum\limits_{{\mathcal{C}}_{1}}+\sum\limits_{{\mathcal{C}}_{2}}\Big)\frac{1}{N}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\\ &\left(q_{0,d_{1}}^{N}(\varphi,0)\right)^{2}\overline{U}_{N}(f_{1}-d_{1})\Big\{\prod\limits_{j=2}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\Big\}\left(q_{f_{k},N}^{N}(0,\psi)\right)^{2},\end{split}$ (5.1)
@@
$\limsup\limits_{N\to\infty}I_{N,\varepsilon}\leq\frac{C}{\log\frac{1}{\varepsilon}}\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2},$ (5.2)
$\limsup\limits_{N\to\infty}J_{N,\varepsilon}\leq C\bigg(\int_{0}^{K_{\varepsilon}\varepsilon}{\mathrm{d}}s\int_{\mathbbm{R}}g_{s}^{2}(x){\mathrm{d}}x+\frac{(\log K_{\varepsilon})^{2}}{\log\frac{1}{\varepsilon}}\bigg)\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2}.$ (5.3)
@@
$\begin{split}\lim\limits_{N\to\infty}\frac{1}{N}&\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}_{k})}\left(q_{0,d_{1}}^{N}(\varphi,0)\right)^{2}\overline{U}_{N}(f_{1}-d_{1})\times\\ &\bigg\{\prod\limits_{j=2}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\bigg\}={\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})\end{split}$ (5.4)
@@
${\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}}):=\!\!\idotsint\limits_{a_{1}\leq b_{1}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{1}),\cdots,a_{k}\leq b_{k}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{k})}\!\!g_{a_{1}}(\varphi,0)^{2}G_{\vartheta}(b_{1}-a_{1}){\mathrm{d}}a_{1}{\mathrm{d}}b_{1}\bigg\{\prod\limits_{j=2}^{k}\frac{G_{\vartheta}(b_{j}-a_{j})}{a_{j}-b_{j-1}}{\mathrm{d}}a_{j}{\mathrm{d}}b_{j}\bigg\},$ (5.5)
@@
Recall the local limit (2.20) for $u(n)$, the asymptotics (2.27) and the uniform bound (2.26) for $\overline{U}_{N}(n)$. The lemma then follows by a Riemann sum approximation, similar to the proof of Proposition 2.3. ${\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})$ is well-defined by the renewal property (6.14) in [13]:
$G_{\vartheta}(t)=\iint_{0<u<\bar{t}\leq v<t}G_{\vartheta}(u)\frac{1}{v-u}G_{\vartheta}(t-v){\mathrm{d}}u{\mathrm{d}}v,\quad\forall t\in(\bar{t},+\infty).$ (5.6)
@@
${\mathcal{I}}_{\varepsilon}^{s,t}({\mathbf{I}}):=\idotsint\limits_{s\leq b_{1}\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{1}),\cdots,a_{k}\leq t\in{\mathcal{T}}_{\varepsilon}(\mathrm{i}_{k})}G_{\vartheta}(b_{1}-s){\mathrm{d}}b_{1}\bigg\{\prod\limits_{j=2}^{k-1}\frac{G_{\vartheta}(b_{j}-a_{j})}{a_{j}-b_{j-1}}{\mathrm{d}}a_{j}{\mathrm{d}}b_{j}\bigg\}\frac{G_{\vartheta}(t-a_{k})}{a_{k}-b_{k-1}}{\mathrm{d}}a_{k},$ (5.7)
@@
$\sum\limits_{k=1}^{\infty}\sum\limits_{\mathrm{j}=:\mathrm{i}_{1}<\mathrm{i}_{2}<\cdots<\mathrm{i}_{k-1}<\mathrm{i}_{k}:=\mathrm{j}^{\prime}}{\mathcal{I}}_{\varepsilon}^{s,t}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=G_{\vartheta}(t-s)$ (5.8)
@@
$\limsup\limits_{N\to\infty}I_{N,\varepsilon}$ $\leq\|\psi\|_{\infty}\sum\limits_{k\geq(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{0<\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}\leq\frac{1}{\varepsilon}}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=:I_{\varepsilon},$ (5.9)
$\limsup\limits_{N\to\infty}J_{N,\varepsilon}$ $\leq\|\psi\|_{\infty}\sum\limits_{k=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in({\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}})^{c}}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})=:J_{\varepsilon}.$ (5.10)
@@
$I_{\varepsilon}\leq\frac{1}{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{{\mathbf{I}}\subset\{1,\cdots,\frac{1}{\varepsilon}\},|{\mathbf{I}}|\geq(\log\frac{1}{\varepsilon})^{2}}|{\mathbf{I}}|{\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}})\leq\frac{1}{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{j=1}^{\frac{1}{\varepsilon}}\sum\limits_{j\in{\mathbf{I}}\subset\{1,\cdots,\frac{1}{\varepsilon}\}}{\mathcal{I}}_{\varepsilon}^{\varphi}({\mathbf{I}}).$